# Creative Preference Optimization

### v1

@@ 6.1 Effect of Injection Weights
While we set all injection weights to 1 for simplicity in our main evaluations, we also study the effect of the different injection values on the performance of models across dimensions. In particular, we vary the injection weights from 00 to $2.0$ with an increment of $0.5$ for all dimensions and report the averaged results across three seed runs in Figure 3. We observe that across most dimensions, an injection weight of $0.5$ yields the greatest performance gains, with further increases resulting in diminishing returns or slight performance degradation. In terms of quality, the injection weight of $1.0$ results in the highest performance. Nevertheless, any weight above 00 consistently outperforms the model without any injection with minimal drop in quality (Appendix Figure 8). We suggest tuning these values depending on the training dataset, underlying task, and the base model for the best performance.
@@ 7 Conclusion
Due to constraints on both computational resources and budget for human studies, we were unable to evaluate CrPO on any languages other than English. Multilingual creativity assessment using generative AI remains a challenging problem and an active area of research Haase et al. (2025). While we believe our data represents a valuable resource for the community, future work will need to test our methods in multilingual settings to ensure multilingual generalization. These compute constraints also prevented us from evaluating CrPO on larger open-weight models, making scaling trends difficult to predict. We retained only samples with full agreement for the creativity score when training our models. While this aligns with best practices for creativity measurement in psychology Cseh and Jeffries (2019), it may also mask genuine sources of rater disagreement that should be modeled. Finally, we acknowledge that, much like other datasets used to align LLMs, the preferences represented by our annotator population likely do not reflect the full range of human preferences, which could bias our models’ generations Yeh et al. (2024). We believe that the large-scale and multilingual nature of our preference data likely makes it one of the most representative creativity datasets currently available, but stress that future work should consider issues of bias and fairness more carefully for LLM creativity assessment.
@@
We also acknowledge the broader debates around the valid use of AI in social-behavioral research Sun et al. (2025) and concerns surrounding AI automation of industries requiring creativity Wilkinson (2023) in which our work is situated. While the over-reliance on AI for creative tasks to the detriment of human welfare is a legitimate concern, AI has also been acknowledged for its potential to enhance human creativity above and beyond what might be possible otherwise de Chantal et al. (2025). Creativity is a vital skill for future knowledge workers to master Forum (2025), and we believe that enhancing the creativity of AI is an important prerequisite for developing AI systems capable of training humans to be more creative.
@@ Acknowledgements
Mete and Lonneke gratefully acknowledge the support of the Swiss National Science Foundation (grant 205121_207437: C - LING). R.E.B. is supported by grants from the US National Science Foundation [DRL-1920653; DRL-240078; DUE-2155070].
@@ Appendix A MuCE Dataset
We then randomly selected items within each task and assigned them to an out-of-distribution item (OOD-i) set. We identified candidate items that corresponded to 5% or less of the responses within a task. Then, for tasks that contained 20 or more total items, we randomly assigned 2 of these items to our OOD-i set. For tasks that contained fewer than 20 total items, we instead randomly assigned 1 of these items to the OOD-i set. Finally, we split the remaining responses in our dataset into training, validation, and out-of-distribution responses (OOD-r) sets according to an 80/10/10 split. We grouped responses into unique combinations of sub-dataset, task, language, item, and rating label, then randomly assigned responses within each combination to each of the sets, ensuring an equal representation of responses associated with each of these variables within the training, validation, and OOD-r sets. Table 1 contains the final dataset statistics for MuCE. Tables 6 and 7 contain the descriptions and data statistics for each task in MuCE. Tables 8, 9, 10, 11, and 12 list some example prompts and low-rated and high-rated responses for each task from MuCE.
@@ Appendix B SFT and Preference Datasets
Past work has shown that data quality is one of the main factors behind preference model performance Liu et al. (2024); Deng et al. (2025); Wang et al. (2024a). In particular, the margin in the score (i.e. reward margin) between the preferred and rejected response may influence the performance of the model, since training pairs with smaller margins are likely to contain annotation noise and be more difficult to learn. We experiment with different reward margins and choose a margin of $5$ for the final experiments as it showed a balance between mitigating annotator noise and creating a dataset with nuanced preferences. Additionally, to ensure a high-quality preference dataset, first we filter the base MuCE dataset and select only the samples that have a full agreement from all annotators. Then we filter out all samples that have a rating below $20$ and limit the number of pairings between samples to $10$. This results in a final preference training dataset of $42,058$ samples (MuCE-Pref). We also create a high-quality instruction-tuning dataset from MuCE-Pref by pairing the prompts with all preferred responses that have a rating above $30$ resulting in a dataset of $5,275$ samples (MuCE-SFT). Tables 2 and 3 contain the statistics for these datasets.
@@ Appendix C Training
SFT model is trained on the MuCE-SFT dataset for a single epoch with a batch size of $2$ per GPU using a gradient accumulation size of $4$ and context size of $1024$. We use a cosine scheduler with a half-cycle warmup and maximum learning rate of $53e-$ 5. Final model achieves $85\%$ mean token accuracy on the validation set.
DPO and CrPO models are trained using the SFT model as a base on our MuCE-Pref dataset for a single epoch with a batch size of $8$ per GPU using a gradient accumulation size of $8$ and context size of $1024$. We use a linear scheduler with a learning rate of $65e-$ 6. All final models achieve over $82\%$ reward accuracy on the validation set.
Figure 8: Effect of injection weights for each dimension on the quality score. Results are averaged across three seed runs.
@@ D.2 NoveltyBench Evaluation
To compute quality, they consider a model of user behavior that describes how users interact with and consume language model generations. They assume that the user has a patience level $p\in[0,1]$: after observing each additional generation, they have a probability $p$ of requesting an additional generation from the language model and observing the next generation, and a probability $1-p$ of stopping interacting with the model. Then they compute the quality of a sequence of generations as the cumulative utility:
@@
Design Solutions Produce solutions to real-world design problems DiStefano et al. (2025) 20 10,818

### v2

@@ 6 Results
We also compare our approach to creative prompting (Brainstorm & Select) and decoding (min-p sampling) strategies which we detail in Appendix D.3 and results are reported in Appendix Tables 2 and 6. We can see that while the creative prompting and decoding strategies improve baseline performance, our models with the standard prompting and decoding approaches still beat them across all dimensions with minimal drop in quality.
@@ 6.1 Effect of Injection Weights
While we set all injection weights to 1 for simplicity in our main evaluations, we also study the effect of the different injection values on the performance of models across dimensions. In particular, we vary the injection weights from 0 to $2.0$ with an increment of $0.5$ for all dimensions and report the averaged results across three seed runs in Figure 3. We observe that across most dimensions, an injection weight of $0.5$ yields the greatest performance gains, with further increases resulting in diminishing returns or slight performance degradation. In terms of quality, the injection weight of $1.0$ results in the highest performance. To gain further insights into the inherent interactions between creativity dimensions, for each dimension, we also report the performance of a model optimized for a given dimension on all the other dimensions and across increasing values of injection weights. Results for this analysis can be found in Appendix Figures 8, 9, 10. Results show that in general, the interaction between different creativity dimensions is complex with some discernible patterns. One clear trend is that the increase in novelty, diversity, and surprise injections results in a quality drop, though not significant. Interestingly, we observe that while an increase in novelty injection follows with overall increase in surprise performance, the opposite does not seem to be true and in fact, more surprise injection seems to cause a drop in novelty. On the other hand, novelty and diversity dimensions seem to have a mutual positive correlation. Finally, we observe no clear relationship between the dimensions of surprise and diversity. In general, we suggest tuning the injection weights depending on the training dataset, underlying task, and the base model for the best performance.
@@ 6.3 NoveltyBench Evaluation
7 Discussion
In this work, we propose a modular preference alignment framework that uses signals from different dimensions of creativity to promote overall output creativity improvement. We show that small LLMs trained using our framework on a psychologically-valid datasets of human creativity outperform strong baselines such as GPT-4o on both automated and human evaluations. Our formulation builds on well-established dimensions from creativity research—specifically novelty, surprise, diversity, and quality—which are frequently identified as core components in both cognitive psychology Modirshanechi et al. (2022); Grace and Maher (2016); Barron (1955) and computational creativity literature Runco and Jaeger (2012); Simonton (2012); Boden (2004). We also employ creativity metrics that provide measurable signals aligning with key cognitive theories and enable practical optimization within language models. These signals have also been shown to provide a good feedback mechanism and increase human creativity in several tasks De Chantal and Organisciak (2023); de Chantal et al. (2025b). Similarly, our work shows that these signals can be used as an implicit feedback mechanism in the learning process to improve language model creativity. Additionally, our framework is well-suited for downstream creative tasks where certain dimensions of creativity matter more than others. In creative writing, novelty and diversity ensure fresh, engaging content, while surprise drives plot twists. In advertising, novelty and surprise capture audience attention in crowded markets. For brainstorming, diversity is key to exploring a wide idea space. By enabling models to be trained with customizable mixes of creativity dimensions, our framework adapts to the needs of each domain.
@@ 8 Conclusion
Due to constraints on both computational resources and budget for human studies, we were unable to evaluate CrPO on any languages other than English. Multilingual creativity assessment using generative AI remains a challenging problem and an active area of research Haase et al. (2025). While we believe our data represents a valuable resource for the community, future work will need to test our methods in multilingual settings to ensure multilingual generalization. These compute constraints also prevented us from evaluating CrPO on larger open-weight models, making scaling trends difficult to predict. We retained only samples with full agreement for the creativity score when training our models. While this aligns with best practices for creativity measurement in psychology Cseh and Jeffries (2019), it may also mask genuine sources of rater disagreement that should be modeled. Moreover, we acknowledge that, much like other datasets used to align LLMs, the preferences represented by our annotator population likely do not reflect the full range of human preferences, which could bias our models’ generations Yeh et al. (2024). We believe that the large-scale and multilingual nature of our collected data likely makes it one of the most representative creativity datasets currently available, but stress that future work should consider issues of bias and fairness more carefully for LLM creativity assessment. Finally, we also acknowledge that creativity is a complex and subjective construct, and the metrics we used may not capture the full richness of human creativity judgments.
@@
We also acknowledge the broader debates around the valid use of AI in social-behavioral research Sun et al. (2025) and concerns surrounding AI automation of industries requiring creativity Wilkinson (2023) in which our work is situated. While the over-reliance on AI for creative tasks to the detriment of human welfare is a legitimate concern, AI has also been acknowledged for its potential to enhance human creativity above and beyond what might be possible otherwise de Chantal et al. (2025a); Loi and van der Plas (2020); Loi et al. (2020). Creativity is a vital skill for future knowledge workers to master Forum (2025), and we believe that enhancing the creativity of AI is an important prerequisite for developing AI systems capable of training humans to be more creative.
@@ Acknowledgements
Mete and Lonneke gratefully acknowledge the support of the Swiss National Science Foundation (grant 205121_207437: C - LING) and Fondazione Aldo e Cele Daccò. Antoine Bosselut gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and a Meta LLM Evaluation Research Grant. Roger E. Beaty is supported by grants from the US National Science Foundation [DRL-1920653; DRL-240078; DUE-2155070].
@@ Appendix A MuCE Dataset
We then randomly selected items within each task and assigned them to an out-of-distribution item (OOD-i) set. We identified candidate items that corresponded to 5% or less of the responses within a task. Then, for tasks that contained 20 or more total items, we randomly assigned 2 of these items to our OOD-i set. For tasks that contained fewer than 20 total items, we instead randomly assigned 1 of these items to the OOD-i set. Finally, we split the remaining responses in our dataset into training, validation, and out-of-distribution responses (OOD-r) sets according to an 80/10/10 split. We grouped responses into unique combinations of sub-dataset, task, language, item, and rating label, then randomly assigned responses within each combination to each of the sets, ensuring an equal representation of responses associated with each of these variables within the training, validation, and OOD-r sets. Table 1 contains the final dataset statistics for MuCE. Tables 8 and 9 contain the descriptions and data statistics for each task in MuCE. Tables 10, 11, 12, 13, and 14 list some example prompts and low-rated and high-rated responses for each task from MuCE.
@@ Appendix B SFT and Preference Datasets
Past work has shown that data quality is one of the main factors behind preference model performance Liu et al. (2024); Deng et al. (2025); Wang et al. (2024a). In particular, the margin in the score (i.e. reward margin) between the preferred and rejected response may influence the performance of the model, since training pairs with smaller margins are likely to contain annotation noise and be more difficult to learn. We experiment with different reward margins and choose a margin of $5$ for the final experiments as it showed a balance between mitigating annotator noise and creating a dataset with nuanced preferences. Additionally, to ensure a high-quality preference dataset, first we filter the base MuCE dataset and select only the samples that have a full agreement from all annotators. Then we filter out all samples that have a rating below $20$ and limit the number of pairings between samples to $10$. This results in a final preference training dataset of $42,058$ samples (MuCE-Pref). We also create a high-quality instruction-tuning dataset from MuCE-Pref by pairing the prompts with all preferred responses that have a rating above $30$ resulting in a dataset of $5,285$ samples (MuCE-SFT). Tables 3 and 4 contain the statistics for these datasets.
Model novelty diversity surprise quality
GPT-4o (std. prompting) $0.04$ $0.18$ $42.63$ $\mathbf{32.94}$
GPT-4o (Brainstorm & Select) $0.05$ $0.28$ $14.26$ $22.01$
CrPO-nov $\mathbf{0.10}$ $0.43$ $143.98$ $27.46$
CrPO-div $\mathbf{0.10}$ $\mathbf{0.49}$ $204.48$ $28.87$
CrPO-sur $\mathbf{0.10}$ $\mathbf{0.49}$ $\mathbf{590.65}$ $17.91$
CrPO-nov-div-sur $\mathbf{0.10}$ $0.47$ $249.31$ $27.07$
CrPO-cre $\mathbf{0.10}$ $0.44$ $203.17$ $28.55$
Table 2: AUT evaluation results comparing a creative prompting strategy such as Brainstorm & Select, to our approach with standard prompting. Best performance is bolded for each metric.
@@ Appendix C Training
SFT model is trained on the MuCE-SFT dataset for a single epoch with a batch size of $2$ per GPU using a gradient accumulation size of $4$ and context size of $1024$. We use a cosine scheduler with a half-cycle warmup and maximum learning rate of 3e53e-5. Final model achieves $85\%$ mean token accuracy on the validation set.
DPO and CrPO models are trained using the SFT model as a base on our MuCE-Pref dataset for a single epoch with a batch size of $8$ per GPU using a gradient accumulation size of $8$ and context size of $1024$. We use a linear scheduler with a learning rate of 5e65e-6. All final models achieve over $82\%$ reward accuracy on the validation set.
Figure 8: Effect of novelty injection weight on all creativity dimensions. Results are averaged across three seed runs.
Figure 9: Effect of diversity injection weight on all creativity dimensions. Results are averaged across three seed runs.
Figure 10: Effect of surprise injection weight on all creativity dimensions. Results are averaged across three seed runs.
@@ D.1 Human Evaluation
Model novelty diversity surprise quality
Llama-3.1-8B $0.05$ $0.21$ $46.28$ $31.59$
Llama-3.1-8B (min-p) $0.06$ $0.22$ $51.42$ $31.31$
CrPO-nov $\mathbf{0.10}$ $0.36$ $71.75$ $32.56$
CrPO-div $0.09$ $\mathbf{0.46}$ $115.73$ $30.42$
CrPO-sur $0.08$ $0.44$ $\mathbf{230.91}$ $27.39$
CrPO-nov-div-sur $0.09$ $0.43$ $120.07$ $29.24$
CrPO-cre $0.09$ $0.39$ $96.92$ $\mathbf{33.51}$
Table 6: Evaluation results comparing a creative decoding strategy min-p sampling, to our approach using Llama-3.1-8B as a base model. Best performance is bolded for each metric.
@@ D.2 NoveltyBench Evaluation
To compute quality, they consider a model of user behavior that describes how users interact with and consume language model generations. They assume that the user has a patience level $p\in[0,1]$: after observing each additional generation, they have a probability $p$ of requesting an additional generation from the language model and observing the next generation, and a probability 1p1-p of stopping interacting with the model. Then they compute the quality of a sequence of generations as the cumulative utility:
@@
D.3 Additional Baselines
We consider additional baseline comparisons using creative prompting and decoding strategies to further show the effectiveness of our approach.
D.3.1 Creative Prompting
We choose an approach based on brainstorming and selecting Summers-Stay et al. (2023) that has shown improvement on creativity scores in the Alternative Uses of Objects Task (AUT). In this method, an iterative prompting strategy is employed by asking the model first to generate a response, then brainstorm about the advantages and drawbacks of its response, and then evaluate the novelty and utility of its response, and propose different uses if needed. We follow the prompts suggested in Summers-Stay et al. (2023) and evaluate one of our best-performing baselines, namely, GPT-4o model on the AUT task with a held-out set of 25 objects where for each object we generate 16 different responses using the same evaluation setup mentioned above. We also evaluate our CrPO models on the same held-out test suite, but with standard single-turn prompting.
D.3.2 Creative Decoding
Our main evaluation setup provides comparisons to popular decoding strategies such as top-p, top-k and temperature sampling that are often used to increase the creativity of language model outputs. However, these approaches are not designed to directly improve output creativity. Hence, to further show the effectiveness of our approach, we consider a recent decoding strategy called min-p sampling that promotes output creativity directly Nguyen et al. (2024). We follow the paper’s recommended setup for best performance (min-p values 0.05, 0.1 with high-temperature values 1.0, 1.5) on creative tasks and evaluate the base model Llama-3.1-8B with this decoding setup on our multi-task evaluation suite.
@@ D.3.2 Creative Decoding
Design Solutions Produce solutions to real-world design problems Luchini et al. (2025); DiStefano et al. (2025) 20 10,818